Calculus II, Spring 2019 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 03

1. Section 10.6: Problems 4, 11, 25, 26, 28, 29, 30, 35, 39, 43, 49, 53.

2. Section 10.7: 7, 11, 15, 19, 23.

Hint: To find the radius of convergence for ∑ a_{n}x^{n}, apply ratio test or root test to ∑ |a_{n}x^{n}|.

3. Find a power series that converges on (1, 3) and diverges otherwise. Do the same for (1, 3], [1, 3) and [1, 3], respectively.